OpenStudy (anonymous):

Use mathematical induction to show 1/2 + 1/4 + ... 1/(2^n) = 1 - 1/(2^(n+1))

OpenStudy (anonymous):

Are you sure you typed it right? Normally with induction, we show that it balances for n=1, then your induction hypothesis ASSUMES it works for n= n. Then you show that it works for the (n + 1)th case. When n= 1, left side becomes 1/2 Right side = 1 - 1/(2^2) = 1-1/4 = 3/4 1/2 doesn't equal 3/4 If it doesn't even work for the first case, I'm not sure how you'll prove it.

OpenStudy (anonymous):

Use mathematical induction to show 1/2 + 1/4 + ... 1/(2^n) = 1 - 1/(2^(n)) *** sorry I typed out half of the original and then half of my notes on the n+1 case

OpenStudy (anonymous):

Okay. Now it makes sense. So you first show that it is true when n=1: 1/2 = 1- (1/2) = 1/2 Now say: Induction Hypothesis: Case where n = k Let us assume that: 1/2 + 1/4 + ... + 1/(2^k) = 1 - 1/(2^k) Now when n = k+1, the left hand side becomes 1/2 + 1/4 +...+1/k + 1/(k+1) 1/2 + 1/4 + ...+1/k = 1-1/(2^k) from our I.H. (Induction Hypothesis) above. Add 1/(k+1) to that. We're almost done.

OpenStudy (anonymous):

I'm going to write it on paper, scan it and attach the file to make it clearer. One minute.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Also would you happen to know the sub-sequential limits for sn = sin(npi/4)

OpenStudy (anonymous):

No sorry.

OpenStudy (anonymous):

Thank you for your help. I appreciate it.

OpenStudy (anonymous):

No problem. Post the other question. Maybe someone else can help.